Jenny and Mark commute to work,travelling in opposite directions.Jenny leaves the house at 8am and averages 35 kph, Mark leaves at 8:15am and averages 40 kph. At what time they will be 40 kilometers apart?

Answer by greenestamps(13215)



Let x be the number of hours after 8am.

Jenny leaves at 8am; Mark leaves 15 minutes (1/4 hour) after 8am.

35x = kilometers Jenny travels
40(x-1/4) = kilometers Mark travels

They are traveling in opposite directions; the distance between them is the sum of the distances each travels.

35x+40(x-1/4)=40
35x+40x-10=40
75x=50
x=2/3

ANSWER: The distance between them will be 40km 2/3 of an hour after 8am, at 8:40 am

CHECK:
35(2/3)+40(2/3-1/4) = 35(2/3)+40(5/12) = 70/3+50/3=120/3 = 40